OpenStudy (anonymous):

factor Comletely: 2x^2-11x+14

OpenStudy (anonymous):

Find two number A and B such that A + B = -11 and A * B = 2 * 14

OpenStudy (anonymous):

 First I can see that there are two negative numbers added together. At the end of this factoring the final product will be a positive number (two negative = a positive) so I know I need two negative signs. What numbers multiplied give 14? 14x1 works! So does 7x2. Let's see what happens with 7x2  it appears these two added give us -11 and positive 14. The really only way to know if this will work is trial and error.